OpenStudy (anonymous):

Integrate this: intf((arcsin(x)/sqrt(1-x^2))dx

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle\int\limits_{~}^{~}\frac{{\rm ArcSin}(x)}{\sqrt{1-x^2}}~dx}\) this ?

OpenStudy (solomonzelman):

what is the derivative of \(\large\color{slate}{\displaystyle{\rm ArcSin}(x)}\) , do you know ?

OpenStudy (anonymous):

yes, the derivatie of arcsin is f '(x) = 1 / sqrt(1 - x 2) right?

OpenStudy (solomonzelman):

yes, 1/sqrt(1-x^2)

OpenStudy (solomonzelman):

So, what u-substitution can you make?

OpenStudy (anonymous):

could you replace arcsin with the derivative of arcsin?

OpenStudy (solomonzelman):

remember, that when dealing with any differentiable function f(x) (case 1) \(\large\color{slate}{\displaystyle\int\limits_{~}^{~}G\left(~ f(x)~\right)\times f'(x)~dx}\) u = f(x) du = f'(x) dx getting you \(\large\color{slate}{\displaystyle\int\limits_{~}^{~}G\left(~ u~\right)~du}\) ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ OR (just) (case 2) \(\large\color{slate}{\displaystyle\int\limits_{~}^{~} f(x)~\times f'(x)~dx}\) u = f(x) du = f'(x) dx getting you \(\large\color{slate}{\displaystyle\int\limits_{~}^{~}u~du}\) ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

OpenStudy (solomonzelman):

So, yes

OpenStudy (solomonzelman):

When you have \(\large\color{slate}{\displaystyle\int\limits_{~}^{~}\frac{{\rm ArcSin}(x)}{\sqrt{1-x^2}}~dx}\) you set \(\large\color{slate}{\displaystyle u={\rm ArcSin}(x)}\)

OpenStudy (solomonzelman):

and the derivative of arcsine \(\large\color{slate}{\displaystyle\frac{1}{\sqrt{1-x^2}}}\), with dx IS REPLACED BY du

OpenStudy (solomonzelman):

(take your time to read if you need to)

OpenStudy (solomonzelman):

(your case, is case 2)

OpenStudy (solomonzelman):

when you can go on, tell me...

OpenStudy (anonymous):

what should I tell you?

OpenStudy (solomonzelman):

well, that you understand everything so far.

OpenStudy (solomonzelman):

if you have questions though, ask

OpenStudy (anonymous):

ok so were at case 2 so we just have to solve it after? so we put u and du together?

OpenStudy (solomonzelman):

yes:)

OpenStudy (solomonzelman):

you have (after u-substitution) \(\large\color{slate}{\displaystyle\int\limits_{~}^{~}u~~du}\)

OpenStudy (solomonzelman):

the denominator and the dx are replaced by du. and arcsin(x) is your u

OpenStudy (solomonzelman):

integrate "u" (apply the power rule). Then, substitute the Arcsin(x) bask for u.

OpenStudy (solomonzelman):

And do not forget the +C as you are done:D

OpenStudy (anonymous):

wait why do we replace the denominator by du?

OpenStudy (solomonzelman):

the denominator, is the derivative of of our "u" (of ArcSin(x) )

OpenStudy (solomonzelman):

so, this is going according to case 2 I showed.

OpenStudy (solomonzelman):

still doesn't make sense?

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle\int\limits_{~}^{~}\frac{ {\rm ArcSin}(x)}{\sqrt{1-x^2}}~dx}\) \(\large\color{gray}{\displaystyle u={\rm ArcSin}(x)}\) \(\large\color{gray}{\displaystyle \frac{du}{dx}=\frac{ 1}{\sqrt{1-x^2}}}\) >>>(therefore) >>> \(\large\color{gray}{\displaystyle du=\frac{dx}{\sqrt{1-x^2}}}\) \(\large\color{slate}{\displaystyle\int\limits_{~}^{~}\color{blue}{\frac{ \color{green}{{\rm ArcSin}(x)}}{\sqrt{1-x^2}}~dx}}\) \(\large\color{slate}{\displaystyle\int\limits_{~}^{~}\color{blue}{ \color{green}{u}~~du}}\)

OpenStudy (solomonzelman):

rate this to make sense on scale 1 (not at all) - 10 (perfectly fine) . for me please.

OpenStudy (anonymous):

6

OpenStudy (solomonzelman):

alright.

OpenStudy (solomonzelman):

what doesn't make sense about it ?

OpenStudy (anonymous):

so I get that it u and du but I'm not sure what to do after.. I know it'll be a fraction but i'm not sure how to get there

OpenStudy (solomonzelman):

well, if you understand everything up to \(\large\color{slate}{\displaystyle\int\limits_{~}^{~}u~du}\) it shouldn't be hard

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle\int\limits_{~}^{~}u~du~~~\Rightarrow~\int\limits_{~}^{~}u^1~du~=~\frac{u^{1\color{red}{+1}}}{\color{red}{(1+1)}}+C=~\frac{~~~~~~?~~~~~~}{?}+C}\)

OpenStudy (solomonzelman):

it is "u" to what power, and divided by what number ?

OpenStudy (anonymous):

oh! so it would be \[\frac{ u^2 }{ 2 }\]

OpenStudy (solomonzelman):

yes, +C

OpenStudy (solomonzelman):

now, remember what your "u" was ?

OpenStudy (anonymous):

oh it's arcsin!

OpenStudy (solomonzelman):

u=ArcSin(x)

OpenStudy (solomonzelman):

so, what will your final answer be?

OpenStudy (solomonzelman):

(substitute back for u)

OpenStudy (anonymous):

\[\frac{ \arcsin(x)^2 }{ 2 }+C\]

OpenStudy (solomonzelman):

yes, \(\large\color{slate}{\displaystyle \frac{\left( {\rm ArcSin}{\color{white}{\LARGE |}}x \right)^2 }{2}+C }\)

OpenStudy (solomonzelman):

if you want, you can verify that by taking the derivative.

OpenStudy (anonymous):

awesome! Thanks so much!!

OpenStudy (solomonzelman):

anytime:)

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle \frac{\color{red}{2}\left( {\rm ArcSin}{\color{white}{\LARGE |}}x \right)^{2\color{red}{-1}} }{2}{\color{blue}{\times\frac{1}{\sqrt{1-x^2}}}}+0 }\)

OpenStudy (solomonzelman):

that is how we differentiate (using the chain rule, in blue).

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle \frac{\color{red}{2}\left( {\rm ArcSin}{\color{white}{\LARGE |}}x \right)^{2\color{red}{-1}} }{2}{\color{blue}{\times\frac{1}{\sqrt{1-x^2}}}}+0 }\) \(\large\color{slate}{\displaystyle \left( {\rm ArcSin}{\color{white}{\LARGE |}}x \right)^{1 }{\color{blue}{\times\frac{1}{\sqrt{1-x^2}}}} }\) so, yes, we did it correctly.

OpenStudy (solomonzelman):

bye!

OpenStudy (solomonzelman):

again, if you ever have a question, then ask.